### v1

@@ Abstract
We study the performance of the linear consensus algorithm on strongly connected graphs using the linear quadratic (LQ) cost as a performance measure. In particular, we derive bounds on the LQ cost by leveraging effective resistance. Our results extend previous analyses—which were limited to reversible cases—to the nonreversible setting. To facilitate this generalization, we introduce novel concepts, termed the back-and-forth path and the pivot node, which serve as effective alternatives to traditional techniques that require reversibility. Moreover, we apply our approach to geometric graphs to estimate the LQ cost without the reversibility assumption. The proposed approach provides a framework that can be adapted to other contexts where reversibility is typically assumed.
@@ 1 Introduction
The analysis of multiagent networks is applied to a variety of subjects, such as multi-robot systems [16], wireless sensor networks [10], and large language models [15]. In information processing of multiagent networks, one of the most simple and fundamental algorithms is linear consensus, also known as average consensus. The linear consensus algorithm is used when each agent has a scalar value, and agents have to match their values to the weighted average of their initial value with limited and decentralized communications. Despite its simplicity, linear consensus algorithm is the basis for many different tasks related to network coordination, such as formation control [20, 25], distributed optimization [29], and cooperative leader following [17, 28].
The performance measurement of linear consensus algorithm has been studied in a various way. Linear consensus network can be understood as a linear discrete-time dynamical system, which is closely related to Markov chain [4], so Markov chain is an important tool for analysis. In fact, the performance of linear consensus algorithm has been analyzed by the speed of convergence using spectral analysis of a transition matrix of a Markov chain [3, 27]. On the other hand, recent studies often adopt approaches from a control theoretic point of view, such as linear quadratic (LQ) cost, which is classical in control. The LQ cost is derived by the summation of $L_{2}$-norm of the difference between the state in each moment and the final state. Moreover, the LQ cost also appears in the error estimation of noisy consensus, a variant of linear consensus with additive random noise. With such background, there are some studies using the LQ cost or similar functions as a performance measure [4, 13, 19]. In these studies, various approaches are employed:
  • Eigenvalue calculation: Eigenvalue are calculated for special graph types, including Cayley graphs, grid graphs, and random geometric graphs [4, 13].
@@
In this paper, we analyze bounds on the LQ cost for nonreversible cases using effective resistance. This generalizes the estimation of the LQ cost for reversible linear consensus networks, as studied in [19], to nonreversible networks. A key advantage of nonreversible Markov chains is its applicability to directed networks. While the reversibility is a useful assumption, it does not hold when communications within the network are directed. Therefore, investigating nonreversible cases is beneficial for understanding linear consensus in directed networks. Discussions on nonreversible Markov are also present in recent studies [5, 26, 6]. In the research of a nonreversible Markov chains, a technique called reversiblization of the transition matrix is used to exploit results of studies on reversible Markov chains [12].
@@
    Although the analogy between the effective resistance and Markov chain does not hold for nonreversible cases, we demonstrate that through the reversiblization of a nonreversible chain, the effective resistance can be applied to a nonreversible linear consensus system.
@@
    We relate the reversiblization and the original nonreversible Markov chain by introducing new notions named the back-and-forth path and pivot node. These concepts serve as alternatives to the notion of $2$-fuzz of an undirected graph [2], which is instrumental in reversible cases, but cannot directly apply to nonreversible cases.
The differences between our approach to upper and lower bounds and those of previous studies are summarized in Table 1. In all methods except our own, reversibility is typically assumed in the analysis of linear consensus, owing to the availability of numerous results pertaining to reversible Markov chains. In addition, we apply our main results to random geometric graphs, which model real-world communication networks and have been extensively studied in the literature (see [11]). For linear consensus on geometric graphs, we derive asymptotic bounds on the LQ cost, yielding results analogous to those in [19] for reversible cases.
Table 1: Approaches on the LQ cost
@@
The remainder of this paper is organized as follows: In Section 2, we introduce the mathematical formulation of linear consensus algorithm, effective resistance, and the relationship between them. In Section 3, we present our main result on the estimation of the LQ cost, along with its proofs. Section 4 demonstrates the application of our result to geometric graphs, and Section 5 concludes the paper.
@@
Let $\mathbb{R}$ be a set of real numbers, $\mathbb{R}^{n}$ be a set of $n$-dimensional real vectors, and $\mathbb{R}^{n\times m}$ be a set of $n\times m$ real matrices. We denote $i$-th entry of $\bm{v}\in\mathbb{R}^{n}$ by $v_{i}$, and $(i,j)$ entry of $A\in\mathbb{R}^{n\times m}$ by $A_{ij}$. Let $\bm{e}_{i}\in\mathbb{R}^{n}$ be a vector whose $i$-th entry is $1$ and other entries are 00, and $\bm{1}\in\mathbb{R}^{n}$ be a vector with all entries equal to $1$. We denote an identity matrix by $I$ and zero matrix by $O$. For a vector $\bm{v}\in\mathbb{R}^{n}$, let $\bm{v}^{\mathsf{T}}$ be a transpose of $\bm{v}$, $\operatorname{span}\{\bm{v}\}$ be a vector space spanned by $\bm{v}$, and $\operatorname{diag}(\bm{v})\in\mathbb{R}^{n\times n}$ be a diagonal matrix satisfying $\operatorname{diag}(\bm{v})_{ii}=v_{i}$ for all $i=1,2,\dots,n$ and $\operatorname{diag}(\bm{v})_{ij}=0$ if $i\neq j$. For a matrix $A$, let $A^{\mathsf{T}}$ be a transpose of $A$, $\operatorname{tr}A$ be a trace of $A$, $\ker A$ be a kernel of $A$, and $A^{\dagger}$ be a Moore-Penrose pseudoinverse of $A$. For two matrices $A$ and $B$, $A\geq B$ means $A_{ij}\geq B_{ij}$ for all $i$ and $j$. We denote a graph with a set of node $V$ and a set of edge $\mathcal{E}$ by $(V,\mathcal{E})$. For an undirected graph, we denote an edge between nodes $u$ and $v$ by $\{u,v\}$, while for a directed graph, we denote an edge from node $u$ to node $v$ by $(u,v)$.
@@ 2.1 Definition of linear consensus algorithm
We model a communication network of $n$ agents labeled $1$ to $n$. Each pair of agents can communicate according to the directed graph $\mathcal{G}=(V,\mathcal{E})$, where $V=\{1,2,\dots,n\}$ represents the label of each agent, and $(i,j)\in\mathcal{E}$ means that $i$ can get information from $j$. We call $\mathcal{G}$ a communication graph. In this paper, the information is represented by a real number.
@@
The constraint that agents can only communicate according to the communication graph $\mathcal{G}$ is reflected to the place of nonzero elements in $P$, which defines the directed graph $\mathcal{G}_{\mathrm{dir}}(P)=(V,\mathcal{E})$, letting $(u,v)\in\mathcal{E}$ if and only if $P_{uv}\neq 0$. Due to the constraint, $\mathcal{G}_{\mathrm{dir}}(P)$ should be a subgraph of the communication graph $\mathcal{G}$. In this paper, we assume $\mathcal{G}_{\mathrm{dir}}(P)=\mathcal{G}$ for simplicity. Furthermore, we define the undirected graph associated with matrix $P$, $\mathcal{G}_{\mathrm{undir}}(P)=(V,\mathcal{E})$, letting $\{u,v\}\in\mathcal{E}$ if and only if $P_{uv}\neq 0$ or $P_{vu}\neq 0$.
Throughout this paper, we assume two properties: $P$ is irreducible, and the diagonal elements of $P$ are positive, where $P$ is irreducible if and only if $\mathcal{G}_{\mathrm{dir}}(P)$ is strongly connected. The positive diagonal assumption leads to the aperiodicity of $P$, which means that the greatest common divisor of the lengths of all cycles in $\mathcal{G}_{\mathrm{dir}}(P)$ is $1$.
By these two assumptions, Perron-Frobenius theorem [23] implies that $P$ has the eigenvalue $1$ with the multiplicity $1$, the corresponding right eigenvector is $\bm{1}$, and corresponding left eigenvector is a strictly positive vector. We denote the right eigenvector by $\bm{\pi}$ normalized so that $\sum_{u}\pi_{u}=1$, and call it the invariant measure of $P$. It also holds that, when $t$ grows to infinity, $P^{t}$ converges to $\bm{1}\bm{\pi}^{\mathsf{T}}$. Because of (1), we obtain $\lim_{t\to\infty}\bm{x}(t)=\bm{1}\bm{\pi}^{\mathsf{T}}\bm{x}(0)$. We also use a diagonal matrix $\varPi:=\operatorname{diag}(\bm{\pi})$.
@@ 2.2 Performance measure of linear consensus algorithm
(2) $J(P):=\sum_{t\geq 0}\|P^{t}-\bm{1}\bm{\pi}^{\mathsf{T}}\|_{\mathrm{F}}^{2}=\frac{1}{n}\operatorname{tr}\left[\sum_{t\geq 0}(I-\bm{\pi}\bm{1}^{\mathsf{T}})(P^{\mathsf{T}})^{t}P^{t}(I-\bm{1}\bm{\pi}^{\mathsf{T}})\right],$
@@
which is obtained by substituting $\|\tilde{\bm{e}}\|^{2}$ in (3) by $\sum_{i}\pi_{i}\tilde{e}_{i}^{2}$. Different from (2), which sums up the errors uniformly, (4) sums up the errors with weights according to $\bm{\pi}$.
@@ 2.3 Effective resistance
We consider a resistor network as an undirected connected graph $\mathcal{G}=(V,\mathcal{E})$ and each edge represents the resistor that connects both nodes of the edge. A resistor network with $n$ nodes is determined by assigning a matrix $C=(C_{ab})\in\mathbb{R}^{n\times n}$ whose element $C_{ab}$ means the conductances of the edge between $a$ and $b$ if $C_{ab}\neq 0$, or that $a$ and $b$ are not connected by an edge if $C_{ab}=0$. We call $C=(C_{ab})$ a conductance matrix if $C$ is a symmetric, nonnegative and irreducible matrix.
@@
Because the current injection and extraction are balanced, we can assume $\bm{i}^{\mathsf{T}}\bm{1}=0$. Under this condition, we can show the following lemma.
@@
Proof 2.2.
Notice that $\operatorname{diag}(C\bm{1})$ is nonsingular and
$\left(I-\operatorname{diag}(C\bm{1})^{-1}C\right)=\operatorname{diag}(C\bm{1})^{-1}L(C).$
Because $\mathcal{G}$ is connected, $\operatorname{diag}(C\bm{1})^{-1}C$ is an irreducible stochastic matrix. Therefore, Perron-Frobenius theorem [23] ensures that $\operatorname{diag}(C\bm{1})^{-1}C$ has an eigenvalue $1$ with multiplicity $1$, so the rank of $I-\operatorname{diag}(C\bm{1})^{-1}C$ is $n-1$, which means that the rank of $L(C)$ is also $n-1$. Because $L(C)\bm{1}=\bm{0}$, $L(C)$ has an eigenvalue 00 with multiplicity $1$. Therefore, $\bm{l}_{1},\dots,\bm{l}_{n-1}$ are linearly independent, where $\bm{l}_{k}^{\mathsf{T}}$ is the $k$-th row of $L(C)$. Taking into account only $i_{1},\dots,i_{n-1}$ and $\bm{l}_{1}^{\mathsf{T}},\dots,\bm{l}_{n-1}^{\mathsf{T}}$ in (5), we can obtain the unique solution $\bm{v}$ for (5) up to the fundamental solution term $\bm{1}$. This $\bm{v}$ automatically satisfies $i_{n}=\bm{l}_{n}^{\mathsf{T}}\bm{v}$, because $i_{n}=-(i_{1}+\dots+i_{n-1})$ and $\bm{l}_{n}=-(\bm{l}_{1}+\dots+\bm{l}_{n-1})$.
Consider the situation that a voltage source is connected between nodes $a,b$, and adjust the voltage so that a unit current flows through the source. In this setting, the currect vector $\bm{i}$ is $\bm{e}_{a}-\bm{e}_{b}$, and $v_{a}-v_{b}$, the potential difference between nodes $a,b$, is unique. Thus, we can define effective resistance between $a,b$:
@@
In our paper, we also consider resistor networks whose resistors have a unit conductance. For such networks, we only have to indicate the unweighted graph $\mathcal{G}$, so we denote the effective resistance between $a,b$ for such networks by $\mathcal{R}_{ab}(\mathcal{G})$, which is the property determined only by the graph topology.
@@
¯$\bar{\mathcal{R}}(C):=\frac{1}{2n^{2}}\sum_{u,v\in V}\mathcal{R}_{uv}(C)$
as the average effective resistance and ¯$\bar{\mathcal{R}}(\mathcal{G})$ is defined using $\mathcal{R}_{uv}(\mathcal{G})$ in the same way. ¯$\bar{\mathcal{R}}(\mathcal{G})$ is also determined only by the graph topology.

### v2

@@ Abstract
We study the performance of the linear consensus algorithm on strongly connected directed graphs using the linear quadratic (LQ) cost as a performance measure. In particular, we derive bounds on the LQ cost by leveraging effective resistance and reversiblization. Our results extend previous analyses—which were limited to reversible cases—to the nonreversible setting. To facilitate this generalization, we introduce novel concepts, termed the back-and-forth path and the pivot node, which serve as effective alternatives to traditional techniques that require reversibility. Moreover, we apply our approach to Cayley graphs and random geometric graphs to estimate the LQ cost without the reversibility assumption. The proposed approach provides a framework that can be adapted to other contexts where reversibility is typically assumed.
@@ 1 Introduction
The analysis of multiagent networks has been applied to a variety of subjects, such as multi-robot systems [17], wireless sensor networks [11], and large language models [16]. One of the simplest and most fundamental algorithms in this context is linear consensus, also known as average consensus. The linear consensus algorithm is used when each agent has a scalar value and have to update its value toward the weighted average of initial values, relying only on limited and decentralized communication. Despite its simplicity, the linear consensus algorithm forms the foundation for many tasks in network coordination, such as formation control [22, 27], distributed optimization [31], and cooperative leader following [18, 30].
The performance measurement of the linear consensus algorithm has been studied in various ways. A linear consensus network can be understood as a linear discrete-time dynamical system closely related to a Markov chain [5], making Markov chain theory a key analytical tool. In fact, the performance of the linear consensus algorithm has been analyzed by the speed of convergence using spectral analysis of the transition matrix of the associated Markov chain [4, 29]. On the other hand, recent studies often adopt approaches from a control-theoretic point of view, such as linear quadratic (LQ) cost, which is classical in control. The LQ cost is defined as the sum of squared $L_{2}$-norm of the difference between the state trajectory and the final state. Moreover, the LQ cost appears in the error estimation of noisy consensus, a variant of linear consensus with additive random noise. Motivated by these connections, several studies have employed the LQ cost or similar functions as performance measures [5, 14, 21]. In these studies, various approaches are employed:
  • Eigenvalue calculation: Eigenvalues are calculated for special graph types, including Cayley graphs, grid graphs, and random geometric graphs [5, 14].
@@
In this paper, we derive both upper and lower bounds on the LQ cost for nonreversible cases using effective resistance. This generalizes the estimation of the LQ cost for reversible linear consensus networks, as studied in [21], to nonreversible networks. A key advantage of nonreversible Markov chains is its applicability to directed networks. While the reversibility is a useful assumption, it does not hold in directed communication scenarios, making the nonreversible setting highly relevant. Thus, recent studies have highlighted the importance of nonreversible Markov chains [6, 28, 7]. A central technique in this line is called reversiblization of the transition matrix, which allows one to exploit results of studies on reversible Markov chains [13].
@@
    Although the analogy between the effective resistance and Markov chain does not hold for nonreversible cases, we demonstrate that through the reversiblization of a nonreversible chain (see Section 3.1), the effective resistance can be applied to a nonreversible linear consensus system.
@@
    We relate the reversiblization and the original nonreversible Markov chain by introducing new notions named the back-and-forth path and pivot node. These concepts serve as alternatives to the notion of $2$-fuzz of an undirected graph [2], which is central in reversible cases, but cannot directly apply to nonreversible cases. These concepts, mainly used in Lemma 3.16, demonstrate how graph-theoretic methods can be adapted to the analysis of nonreversible cases.
The differences between our approach to upper and lower bounds and those of previous studies are summarized in Table 1. In all methods except our own, reversibility is typically assumed in the analysis of linear consensus, owing to the availability of numerous results pertaining to reversible Markov chains.
Furthermore, to clarify the behavior of our bounds, we introduce some specific applications and experiments, including Cayley graphs and random geometric graphs. Cayley graphs are highly symmetric and thus analytically tractable, while random geometric graphs model real-world communication networks. These graphs have been extensively studied in the literature (see [12] and [5]).
Table 1: Approaches to the LQ cost
@@
The remainder of this paper is organized as follows: In Section 2, we introduce the mathematical formulation of linear consensus algorithm, effective resistance, and the relationship between them. In Section 3, we present our main result on the estimation of the LQ cost, along with its proofs. Section 4 demonstrates applications and and numerical experiments, and Section 5 concludes the paper.
@@
Let $\mathbb{R}$ be the set of real numbers, $\mathbb{R}^{n}$ be the set of $n$-dimensional real vectors, and $\mathbb{R}^{n\times m}$ be the set of $n\times m$ real matrices. We denote the $i$-th entry of $\bm{v}\in\mathbb{R}^{n}$ by $v_{i}$, and the $(i,j)$ entry of $A\in\mathbb{R}^{n\times m}$ by $A_{ij}$. Let $\bm{e}_{i}\in\mathbb{R}^{n}$ be the vector whose $i$-th entry is $1$ and other entries are 0, and $\bm{1}\in\mathbb{R}^{n}$ be the vector with all entries equal to $1$. We denote the identity matrix by $I$ and zero matrix by $O$. For a vector $\bm{v}\in\mathbb{R}^{n}$, let $\bm{v}^{\mathsf{T}}$ be the transpose of $\bm{v}$, $\operatorname{span}\{\bm{v}\}$ be a vector space spanned by $\bm{v}$, and $\operatorname{diag}(\bm{v})\in\mathbb{R}^{n\times n}$ be the diagonal matrix satisfying $\operatorname{diag}(\bm{v})_{ii}=v_{i}$ for all $i=1,2,\dots,n$ and $\operatorname{diag}(\bm{v})_{ij}=0$ if $i\neq j$. For a matrix $A$, let $A^{\mathsf{T}}$ be the transpose of $A$, $\operatorname{tr}A$ be the trace of $A$, $\ker A$ be the kernel of $A$, and $A^{\dagger}$ be the Moore-Penrose pseudoinverse of $A$. For two matrices $A$ and $B$, $A\geq B$ means $A_{ij}\geq B_{ij}$ for all $i$ and $j$. We denote a graph with a set of nodes $V$ and a set of edges $\mathcal{E}$ by $(V,\mathcal{E})$. For an undirected graph, we denote an edge between nodes $u$ and $v$ by $\{u,v\}$, while for a directed graph, we denote an edge from node $u$ to node $v$ by $(u,v)$.
We also introduce the following notation for central objects in this paper. The formal definitions are given in Section 2.
Table 2: Central objects in this paper
$P$
A consensus matrix
$P^{*}$
The reversiblization of the consensus matrix $P$
$C$
A conductance matrix
$J(P)$
The linear quadratic cost of the consensus matrix $P$
$J_{\mathrm{w}}(P)$
The weighted version of the linear quadratic cost of
the consensus matrix $P$
$\mathcal{G}_{\mathrm{dir}}(P)$
The associated directed graph of $P$ (the graph where
the edge $(u,v)$ exists if and only if $P_{uv}>0$)
$\mathcal{G}(P)$
The associated undirected graph of $P$ (the graph where
the edge $\{u,v\}$ exists if and only if $P_{uv}>0$ or $P_{vu}>0$
$\mathcal{R}_{uv}(C)$
The effective resistance between the node $u,v$
in the network with the conductance matrix $C$
$\mathcal{R}_{uv}(\mathcal{G}(P))$
The effective resistance between the node $u,v$ in the graph $\mathcal{G}(P)$
(setting the conductance of all edges to 11)
$\bar{\mathcal{R}}(C),\bar{\mathcal{R}}(\mathcal{G}(P))$
The average effective resistance of the network
with the conductance matrix $C$ or the graph $\mathcal{G}(P)$
@@ 2.1 Definition of linear consensus algorithm
We model a communication network of $n$ agents labeled $1$ to $n$. Each pair of agents can communicate according to the directed graph $\mathcal{G}_{\mathrm{dir}}=(V,\mathcal{E})$, where $V=\{1,2,\dots,n\}$ represents the set of agent labels, and $(i,j)\in\mathcal{E}$ means that $i$ can get information from $j$. We call $\mathcal{G}_{\mathrm{dir}}$ a communication graph. In this paper, the information is represented by a real number.
@@
The constraint that agents can only communicate according to the communication graph $\mathcal{G}_{\mathrm{dir}}$ is reflected in the placement of nonzero elements in $P$, that is, the edge $(u,v)$ appears in $\mathcal{G}_{\mathrm{dir}}$ if and only if $P_{uv}\neq 0$. Furthermore, we define the undirected graph associated with matrix $P$, $\mathcal{G}_{\mathrm{undir}}(P)=(V,\mathcal{E})$, letting $\{u,v\}\in\mathcal{E}$ if and only if $P_{uv}\neq 0$ or $P_{vu}\neq 0$. In this paper, we denote $\mathcal{G}_{\mathrm{undir}}=\mathcal{G}$ for simplicity.
Throughout this paper, we assume two properties: $P$ is irreducible, and the diagonal elements of $P$ are positive, where $P$ is irreducible if and only if $\mathcal{G}_{\mathrm{dir}}(P)$ is strongly connected. The positive diagonal assumption may appear to be strong, but it is natural for each agent to use its current state for each iteration, and easy to implement since it does not require any additional communication. In addition, this assumption leads to some useful properties, such as aperiodicity of $P$, which means that the greatest common divisor of the lengths of all cycles in $\mathcal{G}_{\mathrm{dir}}(P)$ is $1$. Aperiodicity is a nice property, it is usual to ensure this property by adding some assumptions. For example, the PageRank algorithm calculates the stationary distribution of a Google matrix, which is arranged from the initial hyperlink matrix to make it aperiodic [20].
By these two assumptions, Perron-Frobenius theorem [25] implies that $P$ has the eigenvalue $1$ with the multiplicity $1$, the corresponding right eigenvector is $\bm{1}$, and corresponding left eigenvector is a strictly positive vector. We denote the left eigenvector by $\bm{\pi}^{\mathsf{T}}$ normalized so that $\sum_{u}\pi_{u}=1$, and call $\bm{\pi}$ the invariant measure of $P$. It also holds that, when $t$ grows to infinity, $P^{t}$ converges to $\bm{1}\bm{\pi}^{\mathsf{T}}$. Because of (1), we obtain $\lim_{t\to\infty}\bm{x}(t)=\bm{1}\bm{\pi}^{\mathsf{T}}\bm{x}(0)$. We also use the diagonal matrix $\varPi:=\operatorname{diag}(\bm{\pi})$.
@@ 2.2 Performance measure of linear consensus algorithm
(2) $J(P):=\frac{1}{n}\sum_{t\geq 0}\|P^{t}-\bm{1}\bm{\pi}^{\mathsf{T}}\|_{\mathrm{F}}^{2}=\frac{1}{n}\operatorname{tr}\left[\sum_{t\geq 0}(I-\bm{\pi}\bm{1}^{\mathsf{T}})(P^{\mathsf{T}})^{t}P^{t}(I-\bm{1}\bm{\pi}^{\mathsf{T}})\right],$
@@
which is obtained by substituting $\|\tilde{\bm{e}}\|^{2}$ in (3) by $\sum_{i}\pi_{i}\tilde{e}_{i}^{2}$. Unlike (2), which sums up the errors uniformly, (4) sums up the errors with weights according to $\bm{\pi}$.
@@ 2.3 Effective resistance
We consider a resistor network as an undirected connected graph $\mathcal{G}=(V,\mathcal{E})$ where each edge represents the resistor that connects both nodes of the edge. A resistor network with $n$ nodes is determined by assigning a matrix $C=(C_{ab})\in\mathbb{R}^{n\times n}$ whose element $C_{ab}$ represents the conductance of the edge between $a$ and $b$ if $C_{ab}\neq 0$, or that $a$ and $b$ are not connected by an edge if $C_{ab}=0$. We call $C=(C_{ab})$ a conductance matrix if $C$ is a symmetric, nonnegative and irreducible matrix.
@@
Because the current injection and extraction are balanced, we can assume $\bm{i}^{\mathsf{T}}\bm{1}=0$. Under this condition, the following lemma is well known fact (e.g., Lemma 6.12 in [3]).
@@
Consider the situation where a voltage source is connected between nodes $a,b$, and adjust the voltage so that a unit current flows through the source. In this setting, the current vector $\bm{i}$ is $\bm{e}_{a}-\bm{e}_{b}$, and $v_{a}-v_{b}$, the potential difference between nodes $a,b$, is unique. Thus, we can define effective resistance between $a,b$:
@@
In our paper, we also consider resistor networks whose resistors have unit conductance. For such networks, we only have to indicate the unweighted graph $\mathcal{G}$, so we denote the effective resistance between $a,b$ for such networks by $\mathcal{R}_{ab}(\mathcal{G})$, which is the property determined only by the graph topology.
@@
$\bar{\mathcal{R}}(C):=\frac{1}{2n^{2}}\sum_{u,v\in V}\mathcal{R}_{uv}(C)$
as the average effective resistance and $\bar{\mathcal{R}}(\mathcal{G})$ is defined using $\mathcal{R}_{uv}(\mathcal{G})$ in the same way. $\bar{\mathcal{R}}(\mathcal{G})$ is also determined only by the graph topology.